# Universal Omni-Language Models

### v2

@@ Abstract
Recent advancements in multimodal large language models (MLLMs) have aimed to integrate and interpret data across diverse modalities. However, the capacity of these models to concurrently process and reason about multiple modalities remains inadequately explored, partly due to the lack of comprehensive modality-wise benchmarks. We introduce OmniBench, a novel benchmark designed to rigorously evaluate models’ ability to recognize, interpret, and reason across visual, acoustic, and textual inputs simultaneously. We define models capable of such tri-modal processing as omni-language models (OLMs). OmniBench is distinguished by high-quality human annotations, ensuring that accurate responses require integrated understanding and reasoning across all three modalities. Our main findings reveal that: i) open-source OLMs exhibit critical limitations in instruction-following and reasoning capabilities within tri-modal contexts; and ii) most baselines models perform poorly (below 50% accuracy) even when provided with alternative textual representations of images or/and audio. These results suggest that the ability to construct a consistent context from text, image, and audio is often overlooked in existing MLLM training paradigms. We advocate for future research to focus on developing more robust tri-modal integration techniques and training strategies to enhance OLM performance across diverse modalities. Codes and live leaderboard could be found at https://m-a-p.ai/OmniBench.
@@ 1 Introduction
The rapid advancement of artificial intelligence has ushered in a new era of multimodal large language models (MLLMs), capable of processing and interpreting diverse data types mainly including images, audio, and text (Li et al., 2024b). These models aim to emulate human-like understanding of the world by integrating information across multiple sensory modalities and learning a comprehensive context from the environment. While significant strides have been made in developing MLLMs that can handle two of the modalities, the ability to concurrently process and reason about the three aforementioned modalities remains a frontier yet to be fully explored.
@@
Figure 1: Example Data from Different Categories. The main contextual information is provided by the corresponding image and audio, while the question and options are expressed with text. Playable demos could be found at the homepage.
The social impact of these MLLMs is far-reaching, providing transformative capabilities for a variety of domains. In healthcare, VLM and ALM can contribute to diagnosing Wu et al. (2023a); Hemdan et al. (2023). The integration of all modalities is expected to significantly improve diagnostic accuracy and patient interaction, making healthcare more accessible and efficient. In urban environments, ALM can contribute to improving safety and traffic management by incorporating urban sound event detection during autonomous driving, such as recognising audio of emergency vehicles to supplementary visual and radar modality Sun et al. (2021). In addition, audio contributes to biodiversity monitoring Terenzi et al. (2021); Liang et al. (2024a) and can be greatly enhanced by MLLM’s ability to analyse both audio and video from a variety of sensors. Finally, it may help robotics or LLM agents to provide better human-computer/robotic interaction (HCI/HRI) service in day-to-day life Liang et al. (2024b); Su et al. (2023).
@@
  • In the setting of using text as the alternatives of audios and images, the open-source VLMs and ALMs show a relatively better results but remain in a preliminary level of capability to understand the given tri-modality context.
  • Compared to the open-source OLMs, the proprietary model seems not suffering from performance drop when it comes to the full evaluation setting of our benchmark.
These results underscore the importance of OmniBench as a tool for identifying areas of improvement and guiding research in multimodal systems. In the following sections, we detail the data collection protocol of OmniBench, present our evaluation results on current state-of-the-art MLLMs, and discuss the implications of our findings for the future of research and development. Through OmniBench, we aim to catalyze advancements in MLLMs, pushing the boundaries of artificial intelligence towards true omni-understanding capabilities.
@@ 2 Related Work
Whisper is an ASR model Radford et al. (2022) and is used for a speech encoder in SALMONN and LTU. BEATs (Chen et al., 2022) is a general audio encoder used in SALMONN, MERT is an acoustic music encoder (Li et al., 2023b) used in Mu-llama and MusiLingo, CLAP is a music-text encoder (Wu et al., 2023c) used in audio-flamingo,
Many studies have been introduced to enable model to perceive complex audio signals and generate text responses based on human instructions. Notable examples in audio perception include SALMONN (Tang et al., 2023), BLSPN (Wang et al., 2023a), Speech-LLaMAN (Wu et al., 2023b), and Qwen-Audio (Chu et al., 2023b), all of which demonstrate promising capabilities in audio-focused dialogues. Regarding the visual modality, the large visual language models, often equipped with pre-trained image encoders (Dosovitskiy, 2020; Touvron et al., 2020; Liu et al., 2021; Radford et al., 2021; Zhai et al., 2023), also makes excellent achievements. BLIP2 (Li et al., 2023a) uses a Q-Former to align the visual knowledge with the textual information during the pre-training phase. LLaVA (Liu et al., 2024b) is pre-trained on the GPT-4 generated multimodal language-image instruction-following data and uses a projection to fuse the vision module and the LM. LLaVA-Next (Liu et al., 2024a) improves the single-image performance with a cost of an increasing number of image tokens. Based on the framework LLaVA mode, QwenVL (Bai et al., 2023), CogVLM (Wang et al., 2023b), and Yi-VL (Young et al., 2024) all achieve significant success through a huge number of pre-training data. However, most existing multi-modal language models only focus on a certain modal and generate textual responses. While some models can process textual, audio, and visual inputs simultaneously, open-source models in this field exhibit relatively average capabilities compared to their closed-source counterparts. In the following context of this literature, we refer to the omni-language models (OLMs) as those capable of processing at least three different modalities of data simultaneously [^1] At the moment, we target the models concurrently processing with image, audio, and text as a starting point since these are the most well-explored modalities in the field..
@@
The vision-language benchmarks aim to test models’ ability to combine visual and language data in tasks like OCR, spatial awareness, multimodal information retrieval (e.g. SciMMIR (Wu et al., 2024a)), and multimodal reasoning skills. MM-Vet (Yu et al., 2023) focuses on visual question answering (VQA), requiring models to interpret visual data and respond to queries. MMBench (Liu et al., 2023b) evaluates models via multiple-choice tasks in both Chinese and English, covering diverse domains. MMStar (Chen et al., 2024a) conducts multi-task evaluations to test multimodal fusion capabilities. MMMU (Yue et al., 2024) and CMMMU (Zhang et al., 2024) assess model performance on complex vision-language tasks, emphasizing sophisticated multimodal reasoning. MMRA Wu et al. (2024b) is designed to evaluate the models’ multi-image relational association capability. In addition, there are several audio-understanding benchmarks. Aishell1 (Bu et al., 2017), Aishell2 (Du et al., 2018), and Librispeech (Panayotov et al., 2015) are designed for automatic speech recognition, while ClothoAQA targets audio QA tasks. For automatic audio captioning and vocal sound classification, researchers have curated Clotho (Drossos et al., 2020) and VocalSound (Gong et al., 2022). However, there is a significant lack of comprehensive understanding benchmarks to assess models’ ability to simultaneously process various modalities, including textual, audio, and visual inputs.
@@ 3 OmniBench
Statistic Action & Activity Story Description Plot Inference Object Identification & Description Contextual & Environmental Identity & Relationship Text & Symbols Count & Quantity Overall
Quantity 251 237 230 211 141 32 25 15 1142
@@
Figure 2: The Data Distribution of OmniBench. Each chart represents the data samples grouped by task types and distinguished by audio types.
3.1 Benchmark Design
Building upon existing multimodal benchmarks, we propose a novel task type categorization in OmniBench that assesses a broad spectrum of reasoning and cognitive abilities. Our taxonomy progresses from fundamental perception (Object Identification & Description) to complex inference (Contextual & Environmental, Identity & Relationship). It incorporates temporal and logical order understanding of events (Action & Activity, Story Description, Plot Inference), spatial awareness (Contextual & Environmental), entity recognition (Object Identification & Description), symbolic processing (Text & Symbols), and quantitative reasoning (Count & Quantity). This comprehensive design evaluates both low-level perceptual skills and high-level cognitive functions, enabling a holistic assessment of multimodal language models’ (MLLMs) capabilities to recognize, describe, integrate information, understand context, and make nuanced inferences. OmniBench comprises 1142 question-answer pairs, with task type distribution, text length, and image and audio characteristics detailed in Table 1. The dataset’s audio content falls into three categories: speech (human vocal communication), sound events (non-speech natural, environmental and mechanical sounds), and music (various compositions and performances). Figure 2 illustrates the distribution of these audio types across task categories. This diverse task set aims to reveal the strengths and weaknesses of existing models in a tri-modal setting, potentially guiding the development of future omni-language models.
Figure 3: OmniBench Annotation Scheme. The example shows flawed data that didn’t pass inspection because the information in audio alone is sufficient to answer.
@@ 3.2 Annotation Protocol
Our annotation scheme is built upon a fundamental principle: the correct answer to each question must require information from both the image and audio components. This ensures that the benchmark effectively evaluates the model’s ability to analyze information across modalities. As shown in Figure 3, we implemented a rigorous annotation pipeline consisting of three stages: initial annotation, human inspection, and model inspection. Data samples that failed to meet our criteria at any stage were returned to annotators for revision, ensuring high-quality, multimodal-dependent samples. Through the whole process, $16$ annotators and $5$ quality inspectors are involved, all of which are full time industrial data annotation employee with higher education backgrounds.
@@
Our quality control process was two-fold, including human inspection round and automatic inspection round assisted by MLLM. First, all annotated instruction-response pairs undergo cross-inspection by human annotators. Inspectors provide detailed reasons for any samples that failed to meet our stringent criteria, allowing for targeted revisions. Samples that pass human inspection are then subjected to a secondary check using a vision-language model LLaVA-1.6-34B (Liu et al., 2024a), where the the automatic quality inspection model is selected by considering trade-off between efficiency and performance. This automated process evaluates each sample under various ablation settings: image and text only, audio transcript and text only, and text only (repeated three times). Samples are only accepted if the model either rejected the task or made mistakes under these limited-information scenarios, confirming the necessity of both visual and auditory information for correct responses. We plot the distribution of the inspection frequency of the passed samples in Figure 4, where we could find that $76\%$ (868) of the passed samples do not require further modification under a well-defined annotation framework and $21.1\%$ of them requiring 1-2 times of revision. During the iterative quality checking, $9.58\%$ ($121$) QA pairs are defined as ”hard to recycle” and dumped after revisions and discussions.
@@ 4 Experiment Settings
Baseline Models.
We select three groups of MLLM baselines according to the modalities available: i. omni-language models: AnyGPT (Zhan et al., 2024), Video-SALMONN (Sun et al., 2024), UnifiedIO2 series (Lu et al., 2024b); ii. vision-language models: InternVL-2 series (Chen et al., 2024b), Qwen2-VL series (team, 2024), Deepseek-VL (Lu et al., 2024a), LLaVA-One-Vision series (Li et al., 2024a), Cambrian series (Tong et al., 2024), Xcomposer2-4KHD (Dong et al., 2024), Idefics2 (Laurençon et al., 2024) as well as the derived Mantis-Idefics2 (Jiang et al., 2024); iii. audio-language models: LTU series (Gong et al., 2023), Mu-LLaMA (Liu et al., 2023a), MusiLingo (Deng et al., 2023), Qwen-Audio series (Chu et al., 2023a), SALMONN-Audio (Sun et al., 2024) and Audio-Flamingo (Kong et al., 2024). We also include the API calls from proprietary models that could support image-text or audio-text inputs, including GPT4-o, Gemini Pro, Reka and Claude-3.5-Sonnet (Achiam et al., 2023; Team et al., 2023; Ormazabal et al., 2024; Anthropic, 2024). We do not categorize them into the group of VLMs, ALMs or OLMs (even not a single model) in our context at the moment since the mechanisms behind these models are not revealed [^2] The authors conclude from an investigation on September 22, 2024..
@@
The main focus of OmniBench is to evaluate how well could the MLLMs understand and reconstruct the context given information from image, audio and text modalities. Setting up questions with four available options for the models, we use accuracy, i.e., the ratio matched letter of the correct option and model response, as the evaluation metric (n.b., the accuracy of a random guess model is $25\%$ under this setting). Additionally, we test the models in an ablation setting of removing one of the image or audio inputs to further reveal a more comprehensive reasoning capability of the baselines and verify the robustness of our benchmark.
@@
For most of the existing MLLMs that only support two input modalities (image-text or audio-text), we build up a simulated evaluation setting allowing us to explore the potential of these models to become omni-languge models in the future. We use the audio transcript annotated by human as the alternative of the audios to enable the evaluation on vision-language models. Regarding the audio-language models, we generate high-quality detailed captions of images automatically with a state-of-the-art VLM, InternVL-2-76B. In such an approximated evaluation setting, models go through the same process of inference and metric calculation as the vanilla one with textual alternatives of images or audios.
Table 2: Overall Omini-Undesratnding Results on Baseline Omni-Language Models. The overall, image-ablated and audio-ablated results on all samples are put at the first three columns.

### v3

@@ Abstract
Recent advancements in multimodal large language models (MLLMs) have aimed to integrate and interpret data across diverse modalities. However, the capacity of these models to concurrently process and reason about multiple modalities remains underexplored, partly due to the lack of comprehensive modality-wise benchmarks. We introduce OmniBench, a novel benchmark designed to rigorously evaluate models’ ability to recognize, interpret, and reason across visual, acoustic, and textual inputs simultaneously. We define language models capable of such tri-modal processing as the omni-language models (OLMs). OmniBench is distinguished by high-quality human annotations, ensuring that accurate responses require integrated understanding and reasoning across all three modalities. Our main findings reveal that: i) open-source OLMs exhibit critical limitations in instruction-following and reasoning capabilities within tri-modal contexts; and ii) most baselines models perform poorly (below 50% accuracy) even when provided with alternative textual representations of images or/and audio. These results suggest that the ability to construct a consistent context from text, image, and audio is often overlooked in existing MLLM training paradigms. To address this gap, we curate an instruction tuning dataset of 84.5K training samples, OmniInstruct, for training OLMs to adapt to multimodal contexts. We advocate for future research to focus on developing more robust tri-modal integration techniques and training strategies to enhance OLM performance across diverse scenarios. Codes, data and live leaderboard could be found at the project page.
@@ 1 Introduction
The rapid advancement of artificial intelligence has ushered in a new era of multimodal large language models (MLLMs), capable of processing and interpreting diverse data types mainly involving images, audio, and text (Li & Lu, 2024). These models aim to emulate human-like understanding of the world by integrating information across multiple sensory modalities and learning a comprehensive context from the environment. While significant strides have been made in developing MLLMs that can handle two of the modalities, the ability to concurrently process and reason about the three aforementioned modalities remains a frontier yet to be fully explored.
The social impact of these MLLMs is far-reaching, providing transformative capabilities for a variety of domains. In healthcare, VLMs and ALMs have contributed to diagnosing (Liu et al., 2023a; Hemdan et al., 2023), and potentially combining three modalities (Meskó, 2023). The integration of all vision, audio and text modalities is expected to significantly improve diagnostic accuracy and patient interaction, making healthcare more accessible and efficient. In urban environments, ALM can contribute to improving safety and traffic management by incorporating urban sound event detection during autonomous driving, such as recognizing audio of emergency vehicles and recognize their types or location with supplementary visual modality (Sun et al., 2021). In addition, audio contributes to biodiversity monitoring (Terenzi et al., 2021; Liang et al., 2024a) and can be greatly enhanced by MLLM’s ability to analyse both audio and video from a variety of sensors. Finally, it may help robotics or LLM agents to provide better human-computer/robotic interaction (HCI/HRI) service in day-to-day life (Liang et al., 2024b; Su et al., 2023).
@@
Figure 1: Example Data from Different Categories. The main contextual information is provided by the corresponding image and audio, while the question and options are expressed with text. Playable audio demos are available at the demo page.
@@
  • Compared to the open-source OLMs, the proprietary models perform better overall but suffer from more accuracy drops when ablating the image or audio input.
  • In the context of using text as an alternative source of information to corresponding audio and images, the open-source VLMs and ALMs show relatively better results but remain in a preliminary level of capability to understand the given tri-modality context.
These results underscore the importance of OmniBench as a tool for identifying areas of improvement and guiding research in multimodal systems. In the following sections, we 1) detail the data collection protocol of OmniBench; 2) present our evaluation results on current state-of-the-art MLLMs; 3) introduce the OmniInstruct dataset for omni-language model supervised fine-tuning; and 4) discuss the implications of our findings for the future of research and development. Through OmniBench, we aim to catalyze advancements in MLLMs, pushing the boundaries of artificial intelligence towards true omni-understanding capabilities.
@@ 2 Related Work
Recent advancements in multimodal large language models (MLLMs) have aimed to integrate and interpret data across diverse modalities. In the audio domain, models like Whisper (Radford et al., 2022), BEATs (Chen et al., 2022), MERT (Li et al., 2023b), and CLAP (Wu et al., 2023b) have been developed as specialized encoders for speech, general audio, acoustic music, and music-text, respectively. These have been incorporated into more comprehensive systems such as SALMONN (Tang et al., 2023), LTU (Gong et al., 2023), Mu-llama, MusiLingo, and Audio-Flamingo. Notable progress in audio perception and instruction-following includes SALMONN (Tang et al., 2023), BLSPN (Wang et al., 2023a), Speech-LLaMAN (Wu et al., 2023a), and Qwen-Audio (Chu et al., 2023b), all demonstrating promising capabilities in audio-focused dialogues. In the visual domain, large visual language models have made significant strides, often leveraging pre-trained image encoders (Dosovitskiy, 2020; Touvron et al., 2020; Liu et al., 2021; Radford et al., 2021; Zhai et al., 2023). Notable examples include BLIP2 (Li et al., 2023a), which uses a Q-Former for visual-textual alignment, LLaVA (Liu et al., 2024b), which employs GPT-4 generated instruction data, and its successor LLaVA-Next (Liu et al., 2024a). Building on the LLaVA framework, models like QwenVL (Bai et al., 2023), CogVLM (Wang et al., 2023b), and Yi-VL (Young et al., 2024) have achieved significant success through extensive pre-training. Despite these advancements, most existing MLLMs focus on a single modality for input processing while generating textual responses. While some models can process textual, audio, and visual inputs simultaneously, open-source models in this field generally exhibit less competitive capabilities compared to their closed-source counterparts. In this context, we define omni-language models (OLMs) as those capable of processing at least three different modalities of data simultaneously [^1] We target the models able to concurrently process image, audio, and text as a starting point since these are the most well-explored modalities in the field, but the ”omni” concept is extendable..
@@
The vision-language benchmarks aim to test models’ ability to combine visual and language data in tasks like OCR, spatial awareness, multimodal information retrieval (e.g. SciMMIR (Wu et al., 2024a)), and multimodal reasoning skills. MM-Vet (Yu et al., 2023) focuses on visual question answering (VQA), requiring models to interpret visual data and respond to queries. MMBench (Liu et al., 2023c) evaluates models via multiple-choice tasks in both Chinese and English, covering diverse domains. MMStar (Chen et al., 2024a) conducts multi-task evaluations to test multimodal fusion capabilities. MMMU (Yue et al., 2024) and CMMMU (Zhang et al., 2024) assess model performance on complex vision-language tasks, emphasizing sophisticated multimodal reasoning. MMRA (Wu et al., 2024b) is designed to evaluate the models’ multi-image relational association capability. In addition, there are several audio-understanding benchmarks. Aishell1 (Bu et al., 2017), Aishell2 (Du et al., 2018), and Librispeech (Panayotov et al., 2015) are designed for automatic speech recognition, while ClothoAQA targets audio QA tasks. For automatic audio captioning and vocal sound classification, researchers have curated Clotho (Drossos et al., 2020) and VocalSound (Gong et al., 2022). However, there is a significant lack of comprehensive understanding benchmarks to assess MLLMs’ ability to simultaneously process complementary information from the textual, audio, and visual inputs.
Audio-Visual Understanding Datasets.
In previous works on audio-visual question answering (AVQA), the focus has predominantly been on identifying visual objects, sounds, and their interrelations to foster multimodal understanding. For instance, the Pano-AVQA dataset (Yun et al., 2021) explores 360-degree panoramic video understanding with 5.4k videos and 51.7k QA pairs. However, it limits its scope to identifying existing objects or locations, omitting questions on causal inference and abstract concepts. Similarly, the Music-AVQA dataset (Li et al., 2022) comprises 9.3k videos, including 1.9k synthesized entries, and 45.9k QA pairs. These videos typically feature simple scenarios of music performances by one or two players or within a music chamber. The questions focus on existing objects, time and location, counting, and relationships but fail to address special symbols like music score images or causal inference. The Music-AVQA-2.0 dataset (Liu et al., 2024c) enhances Music-AVQA by collecting 1,230 manually curated musical ensemble performance videos and 8.1k newly created QA pairs designed to complement and diversify the original dataset, addressing biases in annotation and instrument types. However, it maintains the original types of QA pairs, not expanding into new categories of questions. The AVQA dataset (Yang et al., 2022), contains 57k QA pairs that do not necessarily require integration of both modalities for answering, illustrating its limitation in truly multimodal inquiry. For example, the presence of an auditory signal like a train whistle isn’t essential to deduce an action such as the lowering of a gear lever at a railroad crossing, suggesting that answers could be surmised from visual cues alone. This dataset includes questions on time, location, existing objects, causality, purpose, and counting, yet lacks coverage of actions, symbol concepts, and associations. VALOR dataset (Chen et al., 2023) is an audiovisual-language dataset designed for tri-modality model pre-training, comprising 1.18 M videos sourced and curated from AudioSet (Gemmeke et al., 2017). Recognizing captions derived from ASR or alt-texts fail to adequately align audio-language modalities; VALOR annotates audio-visual content by humans based on AudioSet tags to establish a clear correspondence among 3 modalities. VALOR is good for pre-training but does not include instruction following QA pairs. These gaps in current datasets underscore the need for more comprehensive AVQA datasets and evaluation benchmarks that can challenge and accurately measure a model’s capacity for deep multimodal integration and abstract reasoning, which are critical for advancing multimodal understanding in AI.
@@ 3 OmniBench
Figure 2: The Taxonomy and Proportions of the 1142 Test Samples in OmniBench. The inner circle refers to the three major categories, and the traffic circle refers to the fine-grained task types.
3.1 Benchmark Design
Building on the foundation of existing multimodal benchmarks, our OmniBench introduces a refined taxonomy for task categorization that effectively captures a wide range of cognitive and reasoning abilities. As demonstrate in Figure 2, our framework organizes tasks into three primary categories: (1) (temporal)-spatial entity, which includes Object Identification for recognizing distinct entities and Contextual & Environmental for discerning the setting or backdrop of the events; (2) causal inference, comprised of Story Cause Description to infer narrative drivers, Current Action & Activity to understand ongoing dynamics, and Future Plot and Purpose Inference to anticipate subsequent developments; and (3) abstract concept, involving Identity & Relationship to identify and relate entities, Text & Symbols for symbolic interpretation, and Count & Quantity for numerical reasoning. This taxonomy is designed to evaluate both foundational perceptual skills and complex cognitive processes, thereby providing a comprehensive assessment of multimodal language models’ (MLLMs) abilities to integrate and interpret diverse information sources. OmniBench includes 1142 question-answer pairs, with details on task types, text length, and the characteristics of images and audio presented in Table 1. The audio content of the dataset is categorized into speech, sound events, and music, enriching the diversity of stimuli for evaluating the models’ tri-modal capabilities and aiding in the development of future omni-language models.
@@ 3.1 Benchmark Design
Statistic Causal Inference (Temporal-)Spatial Entity Abstract Concept
Sub-class of QA Current Action & Activity Story Description Plot Inference Object Identification & Description Contextual & Environmental Identity & Relationship Text & Symbols Count & Quantity Overall
Quantity
Total 251 230 237 211 141 32 25 15 1142
Speech 78 182 179 162 104 31 22 13 771
Sound Event 147 27 37 28 25 1 - - 265
Music 26 21 21 21 12 - 3 2 106
@@
Figure 3: OmniBench Annotation Scheme. The annotation example shows flawed data that does not pass inspection because the information in audio alone is sufficient to answer. The audio in the flawed sample will then be sent back to annotators to edit.
@@ 3.2 Annotation Protocol
Our annotation scheme is built upon a fundamental principle: the correct answer to each question must require information from both the image and audio components. This ensures that the benchmark effectively evaluates the model’s ability to analyze information across modalities. As shown in Figure 3, we implemented a rigorous annotation pipeline consisting of three stages: initial annotation, human inspection, and model inspection. Data samples that failed to meet our criteria at any stage were returned to annotators for revision, ensuring high-quality, multimodal-dependent samples. Through the whole process, $16$ annotators and $5$ quality inspectors are involved, all are full-time industrial data annotation employee with higher education backgrounds.
@@
Our quality control process was two-fold, including human inspection round and automatic inspection round assisted by MLLM. First, all annotated instruction-response pairs undergo cross-inspection by human annotators. Inspectors provide detailed reasons for any samples that failed to meet our stringent criteria, allowing for targeted revisions. Samples that pass human inspection are then subjected to a secondary check using a vision-language model LLaVA-1.6-34B (Liu et al., 2024a), where the the automatic quality inspection model is selected by considering trade-off between efficiency and performance. This automated process evaluates each sample under various ablation settings: image and text only, audio transcript and text only, and text only (repeated three times). Samples are only accepted if the model either rejected the task or made mistakes under these limited-information scenarios, confirming the necessity of both visual and auditory information for correct responses. We plot the distribution of the inspection frequency of the passed samples in Figure 4, where we could find that $76\%$ (868) of the passed samples do not require further modification under a well-defined annotation framework and $21.1\%$ of them requiring 1-2 times of revision. During the iterative quality checking, $9.58\%$ ($121$) QA pairs are defined as “hard to recycle” and dumped after revisions and discussions.
3.3 OmniInstruct
To improve the model capability of tri-modal reasoning, we develop the OmniInstruct dataset to facilitate the supervised fine-tuning of models. This dataset leverages the following data sources: the MSRVTT-QA (Xu et al., 2017), AVQA (Yang et al., 2022) and Music-AVQA2.0(Liu et al., 2024c), all of which contain visual, audio and corresponding QA text resources. MSRVTT-QA and AVQA consist of short video clips, typically ranging from 10 to 20 seconds, with minimal scene changes, and music-AVQA 2.0 dataset include 1 minute music performance video. We only adopt the train and validation split of this dataset and regard the whole OmniBench as the test set of the task.
To construct a dataset that aligns with the challenges proposed in OmniBench, we enhanced each question to connect with an audio track and an image extracted from the corresponding video and filter it with VLMs for better quality. Notably, we avoid the first and last five frames of each video to exclude transitional or obscure incomplete scenes that might distort the task’s focus. For MSRVTT-QA train and valid subset, we discard videos without audio tracks and retain a dataset comprised 6,176 videos from the original set that include audio tracks alongside 151.7k QA pairs directly related to these videos. Then we use InternVL-2-76B to filter the questions from the three datasets mentioned above to delete (1) questions that can be answered only with an image, (2) questions irrelevant with image, potentially answerable only with audio, and (3) ambiguous or non-logical questions, where the detailed prompt and statistics could be found at Figure 6 and Table 7, Appendix B. Only 93k data samples remain for training and validation.
This curated dataset is essential for evaluating the nuanced capabilities of multimodal large language models to interpret and integrate multiple types of information, which is a first step towards enhancing reasoning performance and applicability in omni-modality scenarios.
@@ 4 Experiment Settings
Baseline Systems
We select three groups of MLLM baselines according to the modalities available: (i) omni-language models: MIO-Instruct (Wang et al., 2024b), AnyGPT (Zhan et al., 2024), Video-SALMONN (Sun et al., 2024), UnifiedIO2 series (Lu et al., 2024b); (ii) vision-language models: InternVL-2 series (Chen et al., 2024b), Qwen2-VL series (Wang et al., 2024a), Deepseek-VL (Lu et al., 2024a), LLaVA-One-Vision series (Li et al., 2024), Cambrian series (Tong et al., 2024), Xcomposer2-4KHD (Dong et al., 2024), Idefics2 (Laurençon et al., 2024) as well as the derived Mantis-Idefics2 (Jiang et al., 2024); (iii) audio-language models: LTU series (Gong et al., 2023), Mu-LLaMA (Liu et al., 2023b), MusiLingo (Deng et al., 2023), Qwen-Audio series (Chu et al., 2023a), SALMONN-Audio (Sun et al., 2024) and Audio-Flamingo (Kong et al., 2024). We also include the API calls from proprietary models that could support image-text or audio-text inputs, including GPT4-o, Gemini Pro, Reka and Claude-3.5-Sonnet (Achiam et al., 2023; Team et al., 2023; Ormazabal et al., 2024; Anthropic, 2024). We do not conclude them as in the group of VLMs, ALMs or OLMs (even not a single model) in our context at the moment since the mechanisms behind these models are not revealed [^2] The authors conclude from an investigation on September 22, 2024..
@@
The main focus of OmniBench is to evaluate how well could the MLLMs understand and reconstruct the context given information from image ($I$), audio ($A$) and text ($T$) modalities. Setting up questions with four available options for the models, we use accuracy, i.e., the ratio matched letter of the correct option and model response, as the evaluation metric (n.b., the accuracy of a random guess model is $25\%$ under this setting). Additionally, we test the models in an ablation setting of removing one of the image or audio inputs to further reveal a more comprehensive reasoning capability of the baselines and verify the robustness of our benchmark. For baseline systems, please refer to section 4.
@@
For most of the existing MLLMs that only support two input modalities ($(I,T)$ or $(A,T)$), we build up a simulated evaluation setting allowing us to explore the potential of these models to become omni-language models in the future. We use the audio transcript ($A^{\prime}$) annotated by human as the alternative of the audios to enable the evaluation on vision-language models. Regarding the audio-language models, we generate high-quality detailed captions of images ($I^{\prime}$) automatically with a state-of-the-art VLM, InternVL-2-76B. In such an approximated evaluation setting, models go through the same process of inference and metric calculation as the vanilla one with textual alternatives of images or audios.
5 Findings
5.1 Results on Omni-Language Models
Despite overall poor performance, open-source baselines generally exhibit higher accuracy on speech audio, indicating a potential bias towards speech data. In contrast, Gemini-1.5-Pro and Reka-core-20240501, the two available proprietary models evaluated in this tri-modal setting, shows more promising results. Regarding the scores across audio types, the Gemini-1.5-Pro shows a more balanced performances while Reka-core-20240501 showing a lag on modeling the sound events. Besides, Video-Salmonn, developed by Bytedance, and Gemini, developed by Google, provide better results on music subsets compared to their performance on speech and music., potentially due to their large corpus of music videos, though the music ethics of training foundation models are still underdiscussion (Ma et al., 2024). Moreover, the comparison of Gemini-1.5-Pro’s performance across full input context and ablated settings (image-removed and audio-removed) suggests that it effectively leverages information from all modalities to enhance its reasoning capabilities. While it demonstrates superior overall performance and balanced accuracy across audio types compared to open-source alternatives, its accuracy remains below 50%.
Table 2: Overall Omni-Undesratnding Results on Baseline Omni-Language Models. The overall (Image & Audio), image-ablated and audio-ablated results on all samples are provided.